OpenStudy (anonymous):

Solve the equation: square root sign x-6=6

OpenStudy (saifoo.khan):

root (x - 6) = 6 ?

OpenStudy (bahrom7893):

superman is it: sqrt(x-6)=6 or sqrt(x) - 6 = 6?

OpenStudy (bahrom7893):

okay if it's the first one, then: x-6 = 36 x = 42 2nd: sqrtx - 6 = 6 sqrtx = 12 x = 144

OpenStudy (anonymous):

Yess superman it is lol and thanks yall